OpenStudy (anonymous):

With what initial velocity must an object be thrown upward (from ground level) to reach the top of the Washington Monument (approximately 550 ft.)? i have to use integrals to solve and assume the a(t)= -32 feet per second Help working please? i already know the answer

OpenStudy (amistre64):

assuming a(t) = k what is the integral wrt-t ?

OpenStudy (anonymous):

i have to take the integrals of acceleration and velocity to get the answer

OpenStudy (amistre64):

i know, so lets start by integrating a(t) = k, wrt-t

OpenStudy (anonymous):

i've already got acceleration integrated to -32t +C

OpenStudy (amistre64):

good, but lets revise that some int a(t) = k -> v(t) = kt + v , we know at t=0, we want to know the initial velocity "v" now integrate it again v(t) = kt + v , wrt-t to get a function that determines height

OpenStudy (anonymous):

i already have height tho... and i don't do well with letters lol... sorry..but i have tried taking the integral of velocity and i got a position equation but i can't figure out how to use it because i have two variables

OpenStudy (amistre64):

well, show me what you came up with :) im sure your just looking at it a little bit off is all

OpenStudy (anonymous):

yeah i think so..its giving me a headache..

OpenStudy (anonymous):

can i talk to you on skype that way i can just send a picture?

OpenStudy (amistre64):

i dont do skype to follow my idea: int: v(t) = kt + v -> h(t) = kt^2/2 + vt + h , such that when t=0, height=0 giving us h=0 h(t) = kt^2/2 + vt filling in for when h(t)=550, and k=-32 550 = -32t^2/2 + vt

OpenStudy (anonymous):

i can't follow that...i don't understand what your substituting the variables for..

OpenStudy (amistre64):

k is some constant of acceleration, they defined it for us as -32 v is some initial velocity (velocity starts with a "v") that is to be determine. and h is some initial height (height starts with an "h") which we know to be at the ground at a time of zero seconds, so h=0

OpenStudy (amistre64):

the question is really, at what time is the velocity equal to zero?

OpenStudy (anonymous):

oh okay..i'm representing height with s(t)

OpenStudy (anonymous):

yeah..but to find that do i need to find the velocity at 550ft?

OpenStudy (amistre64):

-32t + v = 0 when v = 32t, agreed?

OpenStudy (anonymous):

idk i have never seen that formula

OpenStudy (amistre64):

the velocity at the height of the building is when it has slowed down all the way and is starting to fall again. so the velocity at the top of the building is at zero

OpenStudy (anonymous):

yes i have that written..i meant to say time..sorry..lol

OpenStudy (amistre64):

its your integration of acceleration a(t) integrates to v(t) = -32t + C , where C is the intial velocity that I simply like to call "v" when is the velocity function v(t) equal to zero? well, -32+v = 0 when ... v = 32t: why? (-32t+32t=0)

OpenStudy (amistre64):

so, lets determine our height at v = 32t \[550=-16t^2+vt\] \[550=-16t^2+(32t)t\]

OpenStudy (anonymous):

but i thought you would have to plug in zero to time before you sub. velocity for anything..

OpenStudy (amistre64):

.... might be able to do this in one shot if i keep my head on straight solve v(t) = 0 for t :) -32t + v = 0 when t = v/32 might be more prudent

OpenStudy (amistre64):

\[550=-16t^2+vt\] \[550=-16\frac{v^2}{(-32)^2}+v(\frac{v}{-32})\] and solve for v

OpenStudy (amistre64):

you dont need to sub solve the velcoity equation in order to eliminate one of the variables in the height equation

OpenStudy (amistre64):

either in terms of t = v, or v = t and then do some proper substitutions

OpenStudy (amistre64):

either way, say we go with the first set up; v = 32t 550 = -16t^2 + 32t^2 550 = 16t^2 550/16 = t^2 sqrt(550/16) = t since v(t) = -32t + v 0 = -32(sqrt(550/16)) + v v = 32(sqrt(550/16))

OpenStudy (anonymous):

this is how i have been taught tho..i don't understand exactly how you are getting to -32t + 32t=0

OpenStudy (amistre64):

is it true that we want to determine when the velocity is equal to zero?

OpenStudy (anonymous):

t wouldnt be squared in that second step tho..

OpenStudy (anonymous):

no..im looking for the initial velocity....which will be a big number because of how far the object is traveling

OpenStudy (amistre64):

is it true that we want to know when the velocity function equals out to zero?

OpenStudy (amistre64):

big is relative

OpenStudy (amistre64):

the initial velocity is your +C value of your velocity function

OpenStudy (anonymous):

it will be close to 187.617

OpenStudy (anonymous):

yeah..but i can't plug it in because i don't have it..

OpenStudy (anonymous):

ugh...this is frustrating

OpenStudy (amistre64):

you have 2 equations in 2 unknowns, which can be solved by substitution

OpenStudy (amistre64):

\[550=-16t^2+Ct\\0=-32t+C\] let C = 32t \[550=-16t^2+(32t)t\] \[550=-16t^2+32t^2\] \[550=16t^2\] \[\frac{550}{16}=t^2\] \[\sqrt{\frac{550}{16}}=t\] since C=32t\[C=32\sqrt{\frac{550}{16}}\]

OpenStudy (amistre64):

OR, we could have solved the second equation for "t" instead of "C" \[550=-16t^2+Ct\\0=-32t+C\] let t = -C/32, and sub it into the first equation \[550=-16(-\frac{C}{32})^2+C(\frac{C}{32})\] \[550=\frac{16C^2}{32^2}+\frac{C^2}{32}\] \[550=\frac{16C^2}{32^2}+\frac{32C^2}{32^2}\] \[550=\frac{48}{32^2}C^2\] \[550\frac{32^2}{48}=C^2\] \[\sqrt{550\frac{32^2}{48}}=C\]

OpenStudy (amistre64):

opps, i dropped a negative

OpenStudy (amistre64):

\[550=\frac{-16C^2}{32^2}+\frac{32C^2}{32^2}\] \[550=\frac{16}{32^2}C^2\] thatll be better

OpenStudy (anonymous):

okay..my instructor is telling us to plug in the zero to t in the problem V(t)= -32t + c and then that zero carries through and the problem just tells me that c is initial velocity...but i already know that..so i understand what your doing but i'm not sure if thats what i need to do

OpenStudy (amistre64):

it is what you need to do; define C in terms of t, or define t in terms of C either way ... and sub it in properly to solve

OpenStudy (anonymous):

so don't plug in the zero to t..just replace v and represent v with zero?

OpenStudy (amistre64):

correct, if i read you right we know that at t=0 we have the initial velocity that carries up into the height equation. we want to determine a substitution such that the velocity equation is equal to 0 at some undetermined value of t (or some underdetermined value of C)

OpenStudy (amistre64):

our options are therefore:\[0=-32t+C~:~C=32t\]OR\[0=-32t+C~:~t=\frac{C}{32}\] so the choice is up to you as far as how you want to substitute this into the height function

OpenStudy (amistre64):

personally, i think the C = 32t is nicer to play with :)

OpenStudy (anonymous):

yes it is..i did that..and i understand i think..let me see if i can get the answer

OpenStudy (amistre64):

k, good luck ;)

OpenStudy (anonymous):

i got it=) thank you so much.! sorry if i caused any frustration..lol

OpenStudy (amistre64):

'sok, i needed the practice ;)

OpenStudy (anonymous):

hahaha have a great day;)